Majjor Oil Co. Allocates its maintenance budget to three of its branch offices on the basis(i.e., proportion to) of total floor space. Branch 1 has 9,000 sq.ft., Branch 2 has 20,000 sq.ft., and Branch 3 has 15,000 sq.ft. If $880,000 is to be allocated, how much should each branch receive?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39630)

Total floor space, 44000 square feet
BRANCH       ALLOCATION(dollars)  COMPUTED VALUE REC'D (dollars)
   1         (9/44)880000          180000
   2         (20/44)880000         400000
   3         (15/44)880000         300000

Answer by ikleyn(52878)

.

The long proportion of areas is  9:20:15.


Hence, the allocate amounts are in the same proportion.


It means that Branch 1 gets 9x, Branch 2 gets 20x  and Branch 3 gets 15x,  where x is the common measure.


    9x + 20x + 15x = 880000

    44x            = 880000

      x            = 880000/44 = 20000.


From this point, you can complete the solution on your own.

You received all needed instructions.